OpenStudy (anonymous):

32. If 0 < n < 1, which of the following gives the correct ordering of n,n,and n2? (A) n < n < n2 (D)n

OpenStudy (anonymous):

look at question 32

OpenStudy (anonymous):

here is a trick if you get confused since you are not given numbers, clearly it doesn't matter what they are. so pick one, find the answer, and it will be right no matter what

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i pick \(n=\frac{1}{4}\) so \(n^2=\left(\frac{1}{4}\right)^2=\frac{1}{16}\) and \(\sqrt{n}=\sqrt{\frac{1}{4}}=\frac{1}{2}\)

OpenStudy (anonymous):

now we order them, and whatever answer we get will be the right one this is a good trick to know for standardized tests

OpenStudy (anonymous):

i get \[\frac{1}{16}<\frac{1}{4}<\frac{1}{2}\] so \[n^2<n<\sqrt{n}\] must be correct

OpenStudy (anonymous):

btw i picked \(\frac{1}{4}\) because it was easy to take the square root

OpenStudy (anonymous):

so it's E

OpenStudy (anonymous):

i didn't look at the answers

OpenStudy (anonymous):

ok